how would you do this problem 2x-y=6 3x+4y=-2










Answer by stanbon(75887)

2x-y=6
3x+4y=-2
-----
Modify for elimination:
8x - 4y = 24
3x + 4y = -2
-----
Add and solve for "x":
11x = 22
x = 2
--------
Solve for "y":
2x-y = 6
4-y = 6
y = -2
---------
Cheers,
Stan H.
==================